OpenStudy (anonymous):

Please help! I think the answer is $51,800.00 but I'm not sure. Atlas Chemical Company wanted to keep its recruiting costs for the position of Chief Chemist to less than $100,000. The company spent $15,500 on advertising and $32,700 on interviewing expenses. They will pay 8% commission on selling the selected individual's home. What is the maximum selling price that Atlas would pay 8% commission on and stay within budget?

OpenStudy (madhu.mukherjee.946):

how did you get this answer?

OpenStudy (anonymous):

100,000-15,500-32,700= 51,800 ( took the total amount alloted for the budget and subrtacted the expenses)

OpenStudy (madhu.mukherjee.946):

nope

OpenStudy (anonymous):

s= 647,500 is your answer

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

oh, you divided by .08 .... but why?

OpenStudy (texaschic101):

(15,500 + 32,700) + .08x < 100,000

OpenStudy (anonymous):

it just seems like the number is too big (647500)

OpenStudy (anonymous):

15000+32700+(8%*5)=100000 48200+0.08=100000 -48200 48200 51800 0.08 s=647,500

OpenStudy (anonymous):

OH! ok, i understand now.... thank you guys so much! :)